If the ratio for the sides of a rectangle are 2:5:2:5 and the area of the rectangle is 90, what is the size of each side?
Answer by jim_thompson5910(35256)

A = LW

A = (2x)(5x)

A = 10x^2

90 = 10x^2

90/10 = x^2

9 = x^2

x^2 = 9

x = sqrt(9)

x = 3

------------------

2x = 2*3 = 6

5x = 5*3 = 15

The dimensions of this rectangle is 6 by 15